OpenStudy (lulubj):

Find the product. Write your answer in exponential form. 8^-9 * 8^-6

OpenStudy (trinibutera):

hmmmmmmmmmmmmm

OpenStudy (misty1212):

\[\huge b^m\times b^n=b^{m+n}\] ie add up the exponents

OpenStudy (welshfella):

yea - just add the -9 to -6.

OpenStudy (welshfella):

-9 + -6 = -9-6 = ?

OpenStudy (lulubj):

-15